Dinitrogen tetroxide partially decomposes according to the following equilibrium:

N2O4(g)  2NO2(g)

A 1.000 L flask is charged with 6.00 × 10-2 mol of N2O4. At equilibrium, 3.21 × 10-2 mol of N2O4 remains. Kc for this reaction is ________.

Group of answer choices

0.465

2.42 × 10-2

1.74

10.3

9.70 × 10-2
Answer #1

9 ANS, No Oy - e NO2 Initial 6410-2 moles final 6x1022 moles According to question 6010-2_n = 3.21 x 15² - 2 n = 6410-2 3121equation. According to equilibrium ke = (No22 N₂Oy - (2.79 M10-) 3.2 x 10-2 [Kc = 2.42 x 10-2 Ang